# SynTable: A Synthetic Data Generation Pipeline for Unseen Object Amodal Instance Segmentation of Cluttered Tabletop Scenes

### v2

@@ 6 Conclusion
In conclusion, we present SynTable, a novel synthetic data generation pipeline for generating photorealistic datasets that facilitated amodal instance segmentation of cluttered tabletop scenes. SynTable enables the creation of complex 3D scenes with automatic annotation of diverse metadata, eliminating the need for manual labeling while ensuring dataset quality and accuracy. We demonstrate the effectiveness of the SynTable pipeline by generating a photorealistic amodal instance segmentation dataset and using it to train UOAIS-Net. As a result, UOAIS-Net achieves significantly improved Sim-to-Real transfer performance on the OSD-Amodal dataset, particularly in determining the object occlusion order of objects in a cluttered tabletop scene.
@@ 7 Overview
This supplementary material offers dataset visualization, qualitative results, and additional technical details to support the main paper. Section 8 provides a comprehensive elaboration of the evaluation metrics employed. Section 9 illustrates how occlusion order accuracy is calculated and the validity of the metric. Furthermore, Section 10 delineates the process of generating an occlusion order directed acyclic graph from the occlusion order adjacency matrix to classify objects in three distinct order layers. Lastly, Section 11 showcases some qualitative inference results of UOAIS-Net on the OSD-Amodal dataset. Section 12 presents additional results from our experiments in Section 5 of the main paper. These experiments were conducted by training and evaluating AmodalMRCNN, ORCNN and, ASN on the same datasets as UOAIS-Net. The results observed from the experiments are consistent with our claims in the main paper and further demonstrate the capability of SynTable to improve the performance of a variety of different UOAIS models.
@@ 7.2 Management of the SynTable-Sim Dataset
All the CAD models of the objects used in our SynTable-Sim dataset, as well as the dataset itself, are hosted in the Zenodo open repository, free for all to download. The DOI of our dataset is 10.5281/zenodo.10565517. The dataset can be accessed at https://doi.org/10.5281/zenodo.10565517
@@ 9 Occlusion Order Accuracy $ACC_{oo}$ metric
Correct occlusion order predictions occur when the predicted occlusion relationship for each object matches the ground truth. Incorrect occlusion order predictions can result from erroneous predictions or missing visible mask predictions of object instances. When there are missing predictions, setting the corresponding row and column of the missing object instance in the similarity matrix to 0 penalizes the model for the missing object predictions. The smaller element-wise sum of the similarity matrix leads to a smaller $ACC_{oo}$. This demonstrates the appropriate assignment of penalties by $ACC_{oo}$ to different error types for measuring object occlusion ordering in a scene, highlighting its significance in the context of scene understanding for robotic grasp planning.
@@ 10 Occlusion Order Directed Acyclic Graph (OODAG)
For example, the OOAM generated in Figure 8 shows that $(i,j)=(1,12)$ where $i$ and $j$ are the object indices (the bounding box labels) in the image. This means that object 1 occludes object 12, and a directed edge will point from object 1 to 12. From the generated Directed Occlusion Graph, we can also check if the graph is cyclic or acyclic using graph cyclic detection methods such as Depth First Search (DFS) and Breadth First Search (BFS). Only if the graph has no directed cycles (Directed Acyclic Occlusion Graph) can topological sorting be implemented to find the picking sequence to grasp all objects in the scene safely.
In the generated Occlusion Order graph, we further classify objects in three different order layers - Top, Intermediate, and Bottom. Objects at the top layer represent objects that are not occluded by any other object and can be grasped directly. Objects in the intermediate layers mean that they are occluded but they also occlude other objects. For objects in the bottom layer, they are occluded but they do not occlude other objects.
@@ 11 Qualitative Inference Results of UOAIS-Net on the OSD-Amodal Dataset
After training the UOAIS-Net model [2] on both SynTable-Sim and UOAIS-Sim (tabletop) datasets [2], we present some of our qualitative results in Figure 9. As discussed in the main text of our paper, the UOAIS-Net trained on the SynTable-Sim dataset exhibits superior performance in contrast to the UOAIS-Net trained on the UOAIS-Sim tabletop dataset. This observation is further supported by the inference results presented in Figure 9. Furthermore, as the scene becomes more and more cluttered, the UOAIS-Net model trained on the SynTable-Sim dataset evidently outperforms that of the UOAIS-Net trained on the UOAIS-Sim tabletop dataset. In the context of robotic grasping on cluttered tabletops, foreground masking algorithms can be utilized to filter out the background and out-of-the-table predicted object instances.
@@
12 Additional Quantitative Inference Results on the OSD-Amodal Dataset
We also evaluated the effectiveness of SynTable-Sim across different UOAIS models comprising distinct architectures. Table 5 compares the performance of UOAIS models— Amodal MRCNN, ORCNN, ASN, and UOAIS-Net—on the OSD-Amodal dataset after training on the UOAIS-Sim tabletop dataset and our SynTable-Sim sample dataset. For each model result in our experiments, we used seed 7 for training. Generally, across most metrics, the UOAIS models trained on SynTable-Sim outperform the same models trained on the UOAIS-Sim tabletop dataset. There is also a significant improvement in the results of $ACC_{oo}$ for Amodal MRCNN, ORCNN, and ASN when trained on our SynTable-Sim as compared to the UOAIS-Sim tabletop dataset. This is consistent with the performance trend observed for UOAIS-Net and, therefore, demonstrates that SynTable is an effective tool for generating high-quality datasets that can improve the performance of UOAIS models. A detailed breakdown of the precision P, recall R, and F-measure F, and F@.75 scores for the amodal, invisible, and visible masks are shown in Table 6.
[…]
OV BO F@.75
UOAIS-Sim (Tabletop)
Amodal MRCNN
36.7
26.9
45.7
8.8
4.8
7.7
39.2
54.8
38.7
26.3
32.2
15.6
ORCNN
36.3
25.4
47.0
12.2
6.7
9.0
43.8
59.2
30.5
21.8
29.6
21.5
ASN
40.5
33.6
49.8
17.4
12.1
15.0
47.0
63.2
39.3
31.6
36.8
17.8
UOAIS-Net
49.0
50.3
82.7
42.3
23.9
40.3
68.9
84.0
47.3
50.0
70.6
80.4
SynTable-Sim (Ours)
Amodal MRCNN
74.5
57.5
77.2
41.3
23.5
37.6
69.3
79.4
73.8
57.7
66.1
79.2
ORCNN
74.2
58.2
77.1
44.7
24.3
33.8
72.9
82.2
72.0
58.3
67.7
79.1
ASN
78.2
60.2
75.3
46.4
27.7
35.8
72.6
83.0
78.1
61.8
68.9
80.2
UOAIS-Net
64.4
51.5
84.3
47.3
24.2
47.4
60.0
91.9
65.3
53.7
78.2
87.0
Training Set Method Amodal Mask Invisible Mask Visible Mask
Overlap Boundary F@.75 Overlap Boundary F@.75 Overlap Boundary
F@.75
P R
F
P R
F
P R
F
P R
F
P R
F
P R
F
UOAIS-Sim (Tabletop)
Amodal
27.9 66.7 36.7 22.5 39.8 26.9 45.7 20.2 24.9 8.8 16.4 19.9 4.8 7.7 30.1 60.5 38.7 22.0 37.8 26.3 32.2
MRCNN
ORCNN
26.3
71.1
36.3
19.8
42.4
25.4
47.0
41.5
22.7
12.2
33.9
17.9
6.7
9.0
21.4
63.4
30.5
16.6
38.2
21.8
29.6
ASN
31.7
67.8
40.5
28.6
45.4
33.6
49.8
47.6
23.4
17.4
38.8
20.2
12.1
15.0
32.0
63.5
39.3
28.2
41.5
31.6
36.8
UOAIS-Net
37.2
85.5
49.0
41.1
71.3
50.3
82.7
50.9
54.0
42.3
24.8
41.1
23.9
40.3
35.4
81.6
47.3
41.3
69.3
50.0
70.6
SynTable-Sim (Ours)
Amodal
72.3 81.6 74.5 54.6 64.5 57.5 77.2 54.9 48.0 41.3 30.3 38.8 23.5 37.6 72.1 78.4 73.8 55.1 63.7 57.7 66.1
MRCNN
ORCNN
73.7
80.8
74.2
55.6
64.5
58.2
77.1
61.0
47.1
44.7
31.1
38.6
24.3
33.8
69.8
79.1
72.0
55.7
64.3
58.3
67.7
ASN
78.2
80.3
78.2
57.8
64.9
60.2
75.3
65.2
46.2
46.4
32.9
38.7
27.7
35.8
77.5
80.1
78.1
60.4
65.4
61.8
68.9
UOAIS-Net
53.9
86.3
64.4
40.9
74.6
51.5
84.3
53.0
60.0
47.3
20.5
48.3
24.2
47.4
55.0
86.2
65.3
43.2
75.3
53.7
78.2
[…]
OV BO F@.75
UOAIS-Sim (Tabletop)
Amodal MRCNN
27.1
25.2
23.8
6.7
6.2
3.5
35.8
66.0
29.1
26.2
23.5
19.0
ORCNN
30.9
29.0
28.1
12.5
11.4
8.0
39.9
68.4
31.8
30.2
27.3
23.1
ASN
33.3
34.4
35.3
10.3
9.1
5.0
47.6
72.3
35.0
36.0
34.1
31.6
UOAIS-Net
39.9
40.5
38.6
17.0
15.5
9.6
49.6
74.9
41.6
40.7
35.9
31.6
SynTable-Sim (Ours)
Amodal MRCNN
83.5
76.2
72.5
35.4
31.8
16.4
73.2
80.3
85.7
79.1
71.1
72.8
ORCNN
83.4
76.0
72.2
34.4
29.3
15.3
67.2
73.7
85.3
78.9
70.9
73.0
ASN
83.6
76.9
73.9
38.5
35.1
18.5
74.8
81.5
86.1
80.0
72.8
75.8
UOAIS-Net
83.7
77.5
75.1
40.3
36.7
20.2
75.5
82.0
86.2
80.1
73.3
77.4

### v3

@@ 6 Conclusion
In conclusion, we present SynTable, a novel synthetic data generation pipeline for generating photorealistic datasets that facilitated amodal instance segmentation of cluttered tabletop scenes. SynTable enables the creation of complex 3D scenes with automatic annotation of diverse metadata, eliminating the need for manual labeling while ensuring dataset quality and accuracy. We demonstrate the effectiveness of the SynTable pipeline by generating a photorealistic amodal instance segmentation dataset and using it to train UOAIS-Net. As a result, UOAIS-Net achieves significantly improved Sim-to-Real transfer performance on the OSD-Amodal dataset, particularly in determining the object occlusion order of objects in a cluttered tabletop scene. SynTable advances amodal segmentation for systems that require occlusion-aware perception, such as robotics, augmented reality. By automating annotation of amodal masks and appearance via photorealistic rendering, and scene occlusion order, our pipeline addresses a key bottleneck in training robust vision models with amodal perception capabilities.
@@ 7 Overview
This supplementary material offers dataset visualization, qualitative results, and additional technical details to support the main paper. Section 8 provides additional information about each step in our dataset generation process. Section 9 provides a comprehensive elaboration of the evaluation metrics employed. Section 10 illustrates how occlusion order accuracy is calculated and the validity of the metric. Furthermore, Section 11 delineates the process of generating an occlusion order directed acyclic graph from the occlusion order adjacency matrix to classify objects in three distinct order layers. Lastly, Section 12 showcases some qualitative inference results of UOAIS-Net on the OSD-Amodal dataset.
@@ 7.2 Management of the SynTable-Sim Dataset
The source code for our work is available at https://github.com/ngzhili/SynTable. All the CAD models of the objects used in our SynTable-Sim dataset, as well as the dataset itself, are hosted in the Zenodo open repository, free for all to download. The DOI of our dataset is 10.5281/zenodo.10565517. The dataset can be accessed at https://doi.org/10.5281/zenodo.10565517
8 Additional Details About the Dataset Generation Process
8.1 Preparing Each Scene
8.2 Physical Simulation of Each Scene
8.3 Sampling of Camera Viewpoints
To capture annotations for each scene from multiple viewpoints, we enhance the approach by Gilles et al. [18]—which only uses fixed viewpoint positions—by introducing a feature that captures $V$ number of viewpoints at random positions within two concentric hemispheres, as illustrated in Figure 5. $V$ can be set by the user. The radii of the two concentric hemispheres are uniformly sampled within the range $r_{view\_lower}$ m to $r_{view\_upper}$ m, where $r_{view\_lower}$ and $r_{view\_upper}$ are defined in Equations 2 and 3. Users may also set fixed values for $r_{view\_lower}$ and $r_{view\_upper}$ should they wish to do so.
$r_{view\_lower}=max\left(\frac{w}{2},\frac{l}{2}\right)$ (2)
$r_{view\_upper}=1.7\times r_{view\_lower}$ (3)
The hemisphere’s spherical coordinates are parameterized using three variables $r_{view}$, $u$, and $v$. To generate the camera coordinates in the world frame, we first obtain the radius of the hemisphere $r_{view}$ by uniform sampling between $r_{view\_lower}$ and $r_{view\_upper}$. Next, we uniformly sample $u,v$in [0,1], then substitute all the sampled values into Equations 4, 5 and 6 to compute the cartesian coordinates of the camera.
$x=r_{view}\sin(\arccos(1-v))\cos(2\pi u)$ (4)
$y=r_{view}\sin(\arccos(1-v))\sin(2\pi u)$ (5)
$z=r_{view}\cos(\arccos(1-v))$ (6)
8.4 Sampling of Lighting Conditions
$r_{light\_lower}=r_{view\_upper}+0.1m$ (7)
$r_{light\_upper}=r_{light\_lower}+1m\vspace{-2mm}$ (8)
8.5 Saving of Ground Truth Annotations
We saved the RGB and depth images as PNG images. The OOAM of the objects in each image is saved as a NumPy file. The amodal, visible, and occlusion masks are saved as Run-length Encoding (RLE) in COCO JSON format to optimize disk space used by the generated datasets. We also recorded each object’s visible bounding box, image ID, and object name in the generated COCO JSON file.
Figure 7: Histogram of occlusion rate for UOAIS-Sim tabletop, SynTable-Sim and OSD-Amodal datasets
Figure 8: Histogram for number of regions per connected component (connected component size) for UOAIS-Sim tabletop, SynTable-Sim and OSD-Amodal datasets
@@ 10 Occlusion Order Accuracy $ACC_{oo}$ metric
Correct occlusion order predictions occur when the predicted occlusion relationship for each object matches the ground truth. Incorrect occlusion order predictions can result from erroneous predictions or missing visible mask predictions of object instances. When there are missing predictions, setting the corresponding row and column of the missing object instance in the similarity matrix to 0 penalizes the model for the missing object predictions. The smaller element-wise sum of the similarity matrix leads to a smaller $ACC_{oo}$. This demonstrates the appropriate assignment of penalties by $ACC_{oo}$ to different error types for measuring object occlusion ordering in a scene.
@@ 11 Occlusion Order Directed Acyclic Graph (OODAG)
For example, the OOAM generated in Figure 10 shows that $(i,j)=(1,12)$ where $i$ and $j$ are the object indices (the bounding box labels) in the image. This means that object 1 occludes object 12, and a directed edge will point from object 1 to 12. From the generated Directed Occlusion Graph, we can also check if the graph is cyclic or acyclic using graph cyclic detection methods such as Depth First Search (DFS) and Breadth First Search (BFS). Only if the graph has no directed cycles (Directed Acyclic Occlusion Graph) can topological sorting be implemented.
In the generated Occlusion Order graph, we further classify objects in three different order layers - Top, Intermediate, and Bottom. Objects at the top layer represent objects that are not occluded by any other object. Objects in the intermediate layers mean that they are occluded but they also occlude other objects. For objects in the bottom layer, they are occluded but they do not occlude other objects.
@@ 12 Qualitative Inference Results of UOAIS-Net on the OSD-Amodal Dataset
After training the UOAIS-Net model [2] on both SynTable-Sim and UOAIS-Sim (tabletop) datasets [2], we present some of our qualitative results in Figure 11. As discussed in the main text of our paper, the UOAIS-Net trained on the SynTable-Sim dataset exhibits superior performance in contrast to the UOAIS-Net trained on the UOAIS-Sim tabletop dataset. This observation is further supported by the inference results presented in Figure 11. Furthermore, as the scene becomes more and more cluttered, the UOAIS-Net model trained on the SynTable-Sim dataset evidently outperforms that of the UOAIS-Net trained on the UOAIS-Sim tabletop dataset.